OpenStudy (anonymous):

5] There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $390. How many sophomores bought tickets? 100 90 60 30

OpenStudy (anonymous):

can anybody help me with this question??? plzzz

OpenStudy (anonymous):

f + s =150 2f + 3s = 390 solve the system of equations

OpenStudy (anonymous):

what do i solve for f or s

OpenStudy (anonymous):

same as the other system of equations problems i solved for you you could either use substitution here or just straight up subtract the equations to substitute f = 150 - s 2(150 - s) +3s = 390 then solve for s straight up subtraction (which i think is easier) 2f +2s = 300 - (2f +3s = 390) -------------- -s = -90 s = 90 plug s into f + s = 150 f +90 = 150 f = 60

OpenStudy (anonymous):

so 90 sophomores and 60 freshmen

OpenStudy (anonymous):

o ok so the answer is 60

OpenStudy (anonymous):

no

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

you plug 60 into f , 60 +90=150

OpenStudy (anonymous):

the question is asking how many sophomores there are

OpenStudy (anonymous):

90 sophomores

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanx